# Exploring Modularity of Agentic Systems for Drug Discovery

### v1

@@ Abstract
Large-language models (LLMs) and agentic systems present exciting opportunities to accelerate drug discovery and design. In this study, we critically examine the modularity of LLM-based agentic systems for drug discovery, i.e., whether parts of the agentic system such as the LLM are interchangeable, a topic that has received limited attention in drug discovery applications. We compare the performance of different large language models (LLMs) and the effectiveness of tool-calling agents versus code-generating agents in this domain. Our case study, comparing performance in orchestrating tools for chemistry and drug discovery using an LLM-as-a-judge score, shows that Claude-3.5-Sonnet, Claude-3.7-Sonnet and GPT-4o outperform alternative language models such as Llama-3.1-8B, Llama-3.1-70B, GPT-3.5-Turbo, and Nova-Micro. Although we confirm that code-generating agents outperform the tool-calling ones on average, we show that this is highly question and model dependent. Furthermore, the impact of replacing system prompts is dependent on the specific question asked and the model used, underscoring that—even in this particular domain—one cannot just replace language models without considering prompt re-engineering. Our study highlights the necessity of further research into the modularity of agentic systems to enable the development of stable and scalable solutions for real-world problems.
@@ 1 Introduction
Drug discovery is a complex, multidisciplinary field [31] that has increasingly integrated artificial intelligence (AI) over the past several decades. AI in the pharmaceutical industry has evolved from basic modeling tools in the 1980s to powerful machine learning systems capable of analyzing complex datasets by the end of the 20th century [30]. With continued progress in deep learning and access to vast chemical datasets, AI continuously developed and now significantly enhances the drug discovery and development processes [6]. The next step in AI for drug discovery is anticipated to be the development of research collaborators. These collaborators no longer rely on humans to specify each step and manually transfer the output from one tool to the other, but think along, automate workflows, and ask for input when needed. LLM-based agents that use cheminformatic tools are beginning to bring this to reality [5, 14, 17].
@@
Figure 1: Molecule with its SMILES notation. Brackets correspond to branches, = to a double bond and + and - to positively and negatively charged atoms, respectively.
@@
The progress in large language models (LLMs) in the last few years has been remarkable. Nevertheless, while the reasoning frameworks enhance logical coherence, they lack chemical intuition [4]. In the domain of chemistry, challenges for LLMs remain, such as the use of chemical language [15], especially in the form of SMILES strings (Simplified Molecular Input Line Entry System) [34] strings, which encode molecular structures as text. One of the reasons is that in articles molecules are often described with their common name, IUPAC name and not with SMILES strings, meaning that the training data of most foundational LLM models contains relatively few SMILES strings [19]. Additionally, in SMILES notation, atoms that are close together in the molecular structure are not necessarily close to each other in the text representation (see Figure 1). Also, a molecule can have multiple valid SMILES representations and hydrogen atoms are not explicitly represented. Importantly, models do not have an inherent understanding of SMILES; instead, they treat them as ordinary characters, which adds to the challenge [10]. Navigating the molecular structure and predicting chemical properties is notoriously hard for an LLM [23]. To mitigate the shortcomings of LLMs in chemistry, researchers have begun to augment LLMs with specialized chemistry tools.
@@
Agentic AI – a type of AI system that independently makes decisions and interacts with its surroundings [11] – has already been experimented with in discovery sciences. ChemCrow [24] integrates GPT-4 with external chemical tools using the LangChain framework  [^1] https://github.com/langchain-ai/langchain, enabling it to perform practical chemistry tasks like synthesis planning, molecule generation, and property prediction. PharmAgents [5] is more focused on the pharmaceutical industry, simulating the full drug discovery workflow. Other examples include CoScientist [1] for chemical reaction discovery, DrugAgent for drug discovery tasks such as property prediction [14], assistants for quantum-chemical simulations [12] and Robin for semi-automatic discovery of therapeutic candidates [7]. Although they achieve impressive performance by tailoring the tool orchestration and prompt engineering for this use case, they have not elaborately covered the workflow’s adaptability to new LLMs and tools. Some works claim that the LLM is replaceable, but without studying the effect of doing this [8]. There have been recent benchmarks published on the capabilities of LLMs in the chemical domain [2, 10, 26, 27], yet they mainly focus on the LLM, and, if they use tools, often do not study the use of those tools themselves. Additionally, capabilities are often summarized over a dataset, without considering why some questions are answered more accurately than others, despite research highlighting that the type of question and even the phrasing of the question influence the competitiveness of LLMs [10, 26]. Moreover, LLMs are frequently finetuned to enhance performance [12, 21, 32, 36], which diminishes modularity.
@@
To benefit from the developments both in the field of foundational LLM models and the continuous improvement of cheminformatics tools, these agentic systems should ideally be modular, allowing tools and backbone LLMs to be swapped. However, the prevailing focus on performance has (so far) overshadowed concerns about modularity, leaving this important aspect underexplored.
@@
In this paper, we examine the modularity of LLM-based agentic systems in drug discovery. One of our key contributions is to compare seven different LLMs, revealing that the agentic system’s performance is highly dependent on the underlying LLM. We further compare two common agent types relying, respectively, on tool calling (ToolCallingAgent) and code generation (CodeAgent), also zooming in on the performance on a question-specific level. Finally, we test whether prompt engineering strategies, such as providing clearer examples or domain-specific instructions, yield improvements across different LLMs and questions. On top of this, in support of open-source software, we use the Hugging Face-based SmolAgents framework, demonstrating its application in a drug discovery context for the first time.
@@ 2 Methods
In this section, we cover the agentic framework, the specific types of agents used, and the experimental setup. This includes a description of the smolagent framework and the LLM-as-a-judge used to automatically assess the different agents in our comparison.
@@ 2.1 Agentic Framework
We use the recently released smolagent framework  [^2] https://github.com/huggingface/smolagents, which intends to be more lightweight and flexible compared to alternatives such as the LangChain [25] framework used in ChemCrow [24]. The smolagents framework supports multiple large language models (LLMs) for answering user queries, including those hosted via Azure OpenAI and integrated through LiteLLM 333 https://huggingface.co/docs/smolagents/. Agents in the smolagent framework are multi-step agents that use a ReAct-inspired framework [35] in which actions are carried out by first reasoning and then acting and feeding the execution output to the next step.
The smolagent framework offers different types of agents. The CodeAgent is capable of directly writing and executing code. In this way, the CodeAgent avoids having to change the environment and waiting for the output of a preceding tool call. Previous research [33] shows that the CodeAgent is more flexible and token-efficient than the JSON-based ToolCallingAgent. The ToolCallingAgent is made for more rigid workflows where the tool calls and outcomes are more predictable and less creativity is required. This study sets out to explore the purported benefits of both the smolagents framework and CodeAgents for orchestrating cheminformatics tools.
@@ 2.2 Experimental Setup
Figure 2 presents an overview of the experiments done. First, we compare the performance of seven different LLMs, then we compare the CodeAgent with the ToolCallingAgent, and lastly the influence of changing the system prompt.
@@
To test the tool orchestration capabilities of different agents we formulated 26 industry-representative cheminformatics questions (see Appendix A.1). We consider 17 relatively simple tools (full list in Appendix A.2) that are frequently used in cheminformatics, including functions from the RDkit package [20] and the pubchempy package [^4] https://pubchempy.readthedocs.io/en/latest/ which allows for integration with the PubChem website [18], tools related to physicochemical properties such as the CrippenLogP and molecular weight, but also tools related to converting text into molecular objects and understanding the attributes of molecular objects. These tools are already used to answer industry-relevant questions manually and allow us to formulate multiple valid chains of reasoning ourselves. All LLMs had access to these tools.
In total seven different LLM models were experimented with, including GPT-3.5-Turbo, GPT-4o, Claude-3.5-Sonnet, Claude-3.7-Sonnet, Llama-3.1-8B, Llama-3.1-70B and Nova-Micro (see Appendix A.3 for further details). For all models we experimented with both the ToolCallingAgent and the CodeAgent, yet for the Llama models, the ToolCallingAgent gave an error of not being able to access the tools in our current implementation (see Appendix A.4).
Furthermore, to understand the system prompts’ effect on performance we also tested three versions of the system prompt. The smolagent framework provides a default system prompt that contains an introduction on how the agent should behave and what tools are available. Then it continues with some examples on how to use some of them, followed by a description of the tools and lastly the rules for formatting output and asking the agent to start. For our study, we modified the “default” prompt in two different ways: first, we removed the irrelevant examples that showed how to use fictional tools, to understand if tool usage examples are necessary. This prompt we labeled “cleaner”. Second, we added some context by replacing “You are an expert assistant who can solve any task using tool calls.” with “You are an expert in using tools for computational chemistry who can solve tasks using chemical tools” and adding relevant examples of how to use the specific tools. This prompt we labeled “relevant”.
To score the models we use the the LLM-as-a-judge approach, which means that we use an LLM to score the response of each agent. Following recommendations of a recent survey [9], we include repetitions of each experiment and let the LLM model judge each run independently to generate a score between 0 and 100. A challenge of LLM-as-a-judge is that they often tend to favor responses of a specific length [28]. To mitigate this common problem of length bias we prompted the LLM judge to summarize the answer given by the agent first and then judge the summarized answer (see Appendix A.5). The GPT4o model was used as the judge. The score of the judge ranges between 0 and 100.
@@ 3 Results and Discussions
This section presents the experimental results evaluating the effectiveness of different aspects of the agentic framework across the cheminformatics questions. We start by comparing the performance of the different LLMs and continue by comparing the performance of the CodeAgent and the ToolCallingAgent. Subsequently, we analyze the impact of prompt engineering on task performance and conclude with remarks on noteworthy observations from the experiments.
@@ 3.1 Comparing LLMs
To investigate whether there is a difference in performance between LLMs we consider the score of the CodeAgent with the default prompt of smolagents for the seven different LLMs. For each LLM the 26 questions were asked five times and independently scored by the LLM-as-a-judge system. Figure 3 displays the mean LLM score per model with the standard deviation over the five runs for 26 questions. Clearly, the Claude-3.5-Sonnet, Claude-3.7-Sonnet and GPT-4o have a higher mean score than the other models. A paired one-sided Wilcoxon rank test ($p<0.05$) on the mean value over the 5 repeats per question showed that the Llama models, GPT-3.5-Turbo and the Nova-Micro model perform significantly worse than the best model. Besides the difference in mean performance, the top 3 models show a lower spread in LLM-score across the 26 questions. In conclusion, for this set of questions Claude-3.5-Sonnet, Claude-3.7-Sonnet or GPT-4o perform best. Claude models have previously been shown to perform well in chemistry [26]. While this may suggest that a better domain understanding in the underlying LLM enhances the tool orchestrating capabilities in that domain, a counterexample is the general-purpose biomedical AI agent Biomini for which this was not observed [13].
Looking beyond the average performance, there are a few questions where the on-average-significantly-worse models perform competitively or even better (see Appendix A.6). Examples are questions 2 and 20, which seem to be easy to answer regardless of the LLM involved. The answer to question 2 (“How many hydrogen donors does CC(=O)OC1=CC=CC=C1C(=O)O have?”) is “1 hydrogen donor”. It could be that all models answer this question correctly and use the correct tool calls. Alternatively, however, it is possible that the LLMs guessed the answer and the number 1 is easy to guess, but it could also be that they reasoned from the SMILES instead of using the tools. For question 20 (“How many rotatable bonds does ethene have?”), we see a similar consistent performance. Again, this question has a straightforward answer (“0 rotatable bond”), which could be guessed by the LLM. Ethene is the most basic double-bond molecule, however, meaning that the information for this question is likely included in the training data of the LLM model and no tool calls are required to get this question right. Upon closer inspection, we indeed found that the tools are not always used to answer this question. Furthermore, we can see that for the more complex compound in the form of a SMILES in question 21 (“How many rotatable bonds does CC(C)C1CCC1 have?”) the performance varies more across LLMs and the Claude models no longer answer this question correct 100% of the time. For question 22 (“Based on the number of rotatable bonds, is tabersonine or poloppin a more flexible compound?’), we see a drop in performance, though the chemical property asked about remains the same. Finally, the results on question 8 (“What is the molecular weight of C1OCc2c1ccc3S=CC=Cc23?”) and question 12 (“Does guanine have more hydrogen bond donors or acceptors?”) show other models outperform the Claude models and GPT-4o. This leaves room for further investigations to understand which factors contribute to those specific questions showing different trends across LLMs.
Figure 3: Average LLM score over 26 questions with five repetitions when using the CodeAgent and the default system prompt. Light red indicates that the model performed significantly worse, light green indicates no significant difference and dark green indicates the best model. Bonferroni correction was applied with a significant difference defined as p<0.05.
3.2 ToolCallingAgents vs CodeAgents
Beyond the dependence on the specific LLM, we wanted to investigate the CodeAgents’ performance compared to the ToolCallingAgent. To corroborate that CodeAgents perform better than ToolCallingAgents also in our case study, we compare their LLM scores. We again compare the performance averaged over five runs, quantifying the pairwise difference in score between the two. Figure 4 shows the result as a heatmap indicating whether the CodeAgent (blue) or ToolCallingAgent (red) performs better for that combination of model and question on average. Because of an issue with the particular implementation of the Llama models (see Appendix A.4) available to us, they were excluded from this comparison.
In terms of overall performance, the CodeAgent is better than the ToolCallingAgent in most cases, as there is overall more blue than red. Furthermore, for each of the models, the CodeAgent outperformed the ToolCallingAgent on average (by 2.5 up to 12.9 units). However, a closer look reveals that this is also dependent on the question. For instance, on question 3 (“How many hydrogen acceptors does C1OCc2c1ccc3S=CC=Cc23 have?”), the CodeAgent always performs much better, same for questions 8 and 25 (“What is the molecular weight of C1OCc2c1ccc3S=CC=Cc23?” and “How many nitrogens are there in Wedeloside?”). For question 5 (“What is the molecular weight of caffeine?”), on the other hand, the ToolCallingAgent outperforms the CodeAgent across the models. This is interesting considering that questions 5 and 8 are similar in spirit at first glance, though question 5 might be about more common compounds and thus are more likely to have been represented in the training data of the language models. On top of this, there are several questions where the preference for a ToolCallingAgent versus a CodeAgent depends on the model, such as question 12 (“Does guanine have more hydrogen bond donors or acceptors?”), question 21 (“How many rotatable bonds does CC(C)C1CCC1 have?”) and question 24 (“How many atoms more does oestrogen have than progesterone?”). The exact reason for these differences is still unknown.
@@ 3.2 ToolCallingAgents vs CodeAgents
Figure 4: Difference in LLM score between CodeAgent and ToolCallingAgent for 26 questions and five LLMs. The numbers shown are average over five repeats. The blue areas indicate that the CodeAgent scored higher and the red areas indicate the ToolCallingAgent scored higher.
It could be that the tools that need to be called correlate with a certain LLM preference, but there could also be another aspect of the question impacting the performance, such as the chemical compound that the question is about.

### v2

@@ Abstract
Large-language models (LLMs) and agentic systems present exciting opportunities to accelerate drug discovery. In this study, we examine the modularity of LLM-based agentic systems for drug discovery, i.e., whether parts of the system such as the LLM and type of agent are interchangeable, a topic that has received limited attention in drug discovery. We compare the performance of different LLMs and the effectiveness of tool-calling agents versus code-generating agents. Our case study, comparing performance in orchestrating tools for chemistry and drug discovery using an LLM-as-a-judge score, shows that Claude-3.5-Sonnet, Claude-3.7-Sonnet and GPT-4o outperform alternative language models such as Llama-3.1-8B, Llama-3.1-70B, GPT-3.5-Turbo, and Nova-Micro. Although we confirm that code-generating agents outperform the tool-calling ones on average, we show that this is highly question- and model-dependent. Furthermore, the impact of replacing system prompts is dependent on the question and model, underscoring that even in this particular domain one cannot just replace components of the system without re-engineering. Our study highlights the necessity of further research into the modularity of agentic systems to enable the development of reliable and modular solutions for real-world problems.
@@ 1 Introduction
Drug discovery is a complex, multidisciplinary field [28] that has increasingly integrated artificial intelligence (AI) over the past several decades [27]. The next step in AI for drug discovery is anticipated to be the development of research collaborators, which no longer rely on humans to specify each step and manually transfer the output from one tool to the other, but think independently, automate workflows, and ask for input when needed. LLM-based agents that use cheminformatic tools are starting to bring this to reality [5, 13, 16].
@@
Figure 1: Molecule with its SMILES notation. Brackets correspond to branches, = to a double bond.
@@
The progress in LLMs and their reasoning has been remarkable [34]. While the reasoning frameworks enhance logical coherence, they lack chemical intuition [4]. Chemical language is often challenging for LLMs [14], especially in SMILES (Simplified Molecular Input Line Entry System) [31] strings, which encode molecular structures as text. Reasons for this include that molecules are more often described with their common name or IUPAC name than with SMILES strings, resulting in relatively few SMILES strings in the training data of most foundational LLMs [18]. Additionally, atoms close in the molecular structure are not necessarily close to each other in the text representation, making it more challenging for the LLM to navigate the structure of the molecule (see Figure 1). Importantly, language models do not have an inherent understanding of SMILES and instead treat them as ordinary characters, which adds to the challenge [9]. To mitigate the shortcomings of LLMs, researchers have begun to augment LLMs with specialized chemistry tools.
@@
Agentic AI, a type of AI system that independently makes decisions and interacts with its surroundings [10], has already been experimented with in discovery sciences. Agentic AI could reduce human errors, enable non-technical wet-lab scientists to harness computational tools, and speed up research workflows significantly. For example, ChemCrow [22] integrates GPT-4 with external chemical tools using the LangChain framework [^1] https://github.com/langchain-ai/langchain, enabling it to perform practical chemistry tasks like synthesis planning, molecule generation, and property prediction. PharmAgents [5] is more focused on the pharmaceutical industry, simulating the full drug discovery workflow. Other examples include CoScientist [1] for chemical reaction discovery, DrugAgent for drug discovery tasks such as property prediction [13] and Robin for semi-automatic discovery of therapeutic candidates [6]. Although they achieve impressive performance by tailoring the tool orchestration and prompt engineering for this use case, they have not elaborately covered the workflow’s adaptability to new LLMs and tools. Some works do claim that the LLM is replaceable, yet do not elaborate on the effect of doing this [7]. There have been recent benchmarks published on the capabilities of LLMs in the chemical domain [2, 9, 23, 24], yet they focus on the LLM not agentic AI, and, even if they use tools, often do not study the use of those tools themselves. Additionally, capabilities are often summarized over a dataset, without considering why some questions are answered more accurately than others, despite research highlighting that the type of question and even the phrasing of the question influence the competitiveness of LLMs [9, 23]. Besides, LLMs are frequently finetuned to enhance performance [11, 20, 29, 33], which means that LLMs cannot just be replaced without retraining.
@@
To benefit from the developments both in the field of foundational LLM models and the continuous improvement of cheminformatics tools, these agentic systems should be modular, allowing tools and backbone LLMs to be swapped and prompt improvements to lead to consistent improvement independent of the model. However, the prevailing focus on performance has (so far) overshadowed concerns about modularity.
@@
In this paper, we examine the modularity of LLM-based agentic systems in drug discovery. One of our key contributions is to compare seven LLMs, revealing that the agentic system’s performance is highly dependent on the underlying LLM. We further compare the JSON-based tool calling (ToolCallingAgent) and code generation (CodeAgent) agent type, also zooming in on the performance on a question-specific level. Finally, we test whether prompt engineering strategies, such as providing clearer examples or domain-specific instructions, yield improvements across different LLMs and questions. Additionally, we employ the lightweight, Hugging Face-based smolagents framework and demonstrate its application in a drug discovery context for the first time.
@@ 2 Methods
In this section, we cover the agentic framework, the specific types of agents used, and the experimental setup. This includes a description of the smolagents framework and the LLM-as-a-judge used to automatically assess the different agents in our comparison.
@@ 2.1 Agentic Framework
We use the recently released smolagents framework [^2] https://github.com/huggingface/smolagents, which intends to be more lightweight and flexible compared to alternatives such as LangChain [^3] https://www.langchain.com/. The smolagents framework supports multiple LLMs for answering user queries, including those hosted via Azure OpenAI and integrated through LiteLLM444 https://huggingface.co/docs/smolagents/. Agents in the smolagents framework are multi-step agents that use a ReAct-inspired system framework [32] in which actions are carried out by first reasoning and then acting and feeding the execution output to the next step.
The smolagents framework offers different types of agents. The CodeAgent is capable of directly writing and executing code and in this way avoids having to change the environment and waiting for the output of a preceding tool call. Previous research [30] shows that the CodeAgent is more flexible and token-efficient than the JSON-based ToolCallingAgent. The ToolCallingAgent is suited for more rigid workflows where the tool calls and outcomes are more predictable and less creativity is required. This study sets out to explore the claimed benefits of both the smolagents framework and CodeAgents for orchestrating cheminformatics tools.
@@ 2.2 Experimental Setup
Figure 2 presents an overview of the experiments done. First, we compare the performance of seven different LLMs, then the CodeAgent with the ToolCallingAgent, and lastly the study influence of adapting the system prompt.
@@
To test the tool orchestration capabilities of different agents we formulated 26 industry-representative cheminformatics questions (see Appendix A.1). We consider 17 relatively simple tools (full list in Appendix A.2), including functions from the RDkit package [19] and the pubchempy package [^5] https://pubchempy.readthedocs.io/en/latest/. This allows for integration with the PubChem website [17] and gives us tools related to physicochemical properties such as the CrippenLogP and molecular weight. Importantly, we also include tools related to converting text into molecular objects and understanding the attributes of molecular objects. These tools are already used to answer industry-relevant questions manually and allow us to formulate multiple valid chains of reasoning ourselves. All LLMs had access to these tools.
In total seven different LLM models were experimented with, including GPT-3.5-Turbo, GPT-4o, Claude-3.5-Sonnet, Claude-3.7-Sonnet, Llama-3.1-8B, Llama-3.1-70B and Nova-Micro (see Appendix A.3 for further details). For all models we experimented with both the ToolCallingAgent and the CodeAgent, yet for the Llama models, the ToolCallingAgent gave an error (see Appendix A.4).
Furthermore, to understand the system prompts’ effect on performance we also tested three versions of the system prompt (see Appendix A.5 for the default system prompt of the CodeAgent). We modified the “default” prompt by first removing irrelevant examples of fictional tool usage to create a “clean” prompt, and then added computational chemistry context, replacing the general instructions and including relevant examples to get a “relevant” prompt.
To score the models we use the LLM-as-a-judge approach using GPT4o as the judge. Following recommendations of a recent survey [8], we include repetitions of each experiment and let the LLM model judge each run independently to generate a score between 0 and 100. To mitigate the common problem of length bias [25] we prompted the LLM judge to summarize the answer given by the agent first and then judge the summarized answer (see Appendix A.6).
@@ 3 Results and Discussions
Figure 3: Average with standard deviation LLM score over 26 questions with five repetitions when using the CodeAgent and the default system prompt. Light red indicates a model that performed significantly worse, light green indicates no significant difference and dark green indicates the best model. Bonferroni correction was applied with a significant difference defined as $p<0.05$.
This section presents the experimental results evaluating the effectiveness of different aspects of the agentic framework across the cheminformatics questions. We start by comparing the performance of the different LLMs and continue by studying the performance of the CodeAgent and the ToolCallingAgent. Subsequently, we analyze the impact of prompt engineering on task performance.
@@ 3.1 Comparing LLMs
To investigate the difference in performance between LLMs we evaluated the CodeAgent with the smolagents default system prompt for the seven different LLMs. For each LLM the 26 questions were asked five times and independently scored by the LLM-as-a-judge system. As shown in Figure 3 Claude-3.5-Sonnet, Claude-3.7-Sonnet and GPT-4o achieve a higher mean score than the other models and lower standard deviation. A paired one-sided Wilcoxon rank test ($p<0.05$) on the mean value over the 5 repeats per question showed that the Llama models, GPT-3.5-Turbo and the Nova-Micro model perform significantly worse than the best model. Claude models have previously been shown to perform well in chemistry [23]. While this may suggest that a better domain understanding in the underlying LLM enhances the tool orchestrating capabilities in that domain, a counterexample is the general-purpose biomedical AI agent Biomini for which this was not observed [12].
@@
Figure 4: LLM score difference between CodeAgent and ToolCallingAgent across 26 questions and 5 LLMs (average over 5 runs).
Looking beyond average performance, there are a few questions where models that perform worse overall match or even surpass the top-performing ones (see Appendix A.7). All models answer question 20 (“How many rotatable bonds does ethene have?”), correctly for instance. This question has a straightforward answer (“0 rotatable bond”), which could be guessed. However, since ethene is the most basic double-bond molecule, this information is likely included in the training data of the LLM. Upon closer inspection, we indeed found that the tools are not always used to answer this question. Furthermore, we can see that for the more complex compound in the form of a SMILES in question 21 (“How many rotatable bonds does CC(C)C1CCC1 have?”) the performance varies more across LLMs and top models no longer always answer this question correct. Finally, the results on question 8 (“What is the molecular weight of C1OCc2c1ccc3S=CC=Cc23?”) and question 12 (“Does guanine have more hydrogen bond donors or acceptors?”) show other models outperform the Claude models and GPT-4o. This leaves room for further investigations to understand which factors contribute to those specific questions showing different trends across LLMs.
3.2 ToolCallingAgents versus CodeAgents
Beyond the reliance on the specific LLM, to corroborate that CodeAgents perform better than ToolCallingAgents also in our case study, we compare their LLM scores. Figure 4 shows the result as a heatmap indicating whether the CodeAgent (blue) or ToolCallingAgent (red) performs better for that combination of model and question on average.
In terms of overall performance, the CodeAgent is better than the ToolCallingAgent in most cases, as indicated by a greater prevalence of blue compared to red and for each of the models, the CodeAgent outperformed the ToolCallingAgent on average (by 2.5 up to 12.9 units). However, a closer look reveals that this is also dependent on the question. For instance, on question 3 (“How many hydrogen acceptors does C1OCc2c1ccc3S=CC=Cc23 have?”), the CodeAgent always performs much better, same for questions 8 and 25 (“What is the molecular weight of C1OCc2c1ccc3S=CC=Cc23?” and “How many nitrogens are there in Wedeloside?”). For question 5 (“What is the molecular weight of caffeine?”), on the other hand, the ToolCallingAgent outperforms the CodeAgent across the models. This is interesting considering that questions 5 and 8 are similar in spirit at first glance. On top of this, there are several questions where the preference for a ToolCallingAgent versus a CodeAgent depends on the model, such as question 12 (“Does guanine have more hydrogen bond donors or acceptors?”), question 21 (“How many rotatable bonds does CC(C)C1CCC1 have?”) and question 24 (“How many atoms more does oestrogen have than progesterone?”). It could be that the tools that need to be called correlate with a certain LLM preference, but there could also be another aspect of the question impacting the performance, such as the chemical compound that the question is about. Further research is required to pinpoint the reason for this.